OpenStudy (anonymous):

Part 1: Explain, in complete sentences, which method you would use to solve the following system of equations. (1 point) Part 2: Explain why you chose that method. (1 point) Part 3: Provide the solution to the system. (2 points) x – 3y + 4z = –9 2x + y + 3z = 6 x – 4y + 3z = –9

OpenStudy (anonymous):

@satellite73 @zepdrix @kropot72 @SolomonZelman anyway one of you could possiblt help me with this one?

OpenStudy (anonymous):

@Compassionate @abb0t @nincompoop @Preetha can one of you help please?

OpenStudy (anonymous):

@TTP here is a link to the solutions. they do not match yours. http://www.wolframalpha.com/input/?i=x+%E2%80%93+3y+%2B+4z+%3D+%E2%80%939+and+2x+%2B+y+%2B+3z+%3D+6+and+x+%E2%80%93+4y+%2B+3z+%3D+%E2%80%939

OpenStudy (ttp):

Well I've answered this question twice today and i've got them right! :)

OpenStudy (anonymous):

Ok, im trusting you on this!!! Thank you!!

OpenStudy (ttp):

Your welcome! :)

OpenStudy (anonymous):

@TTP i just read over your work and they don't match up to the question. Maybe the once you answered earlier had different numbers. For example "Subtract equation 2 from equation 1: -5y-z=-18" if i were to subtract equation 2 from equation 1 i would not get that

OpenStudy (ttp):

Trust me! Have you sumbitted it?

OpenStudy (anonymous):

@TTP no, because the work doesn't match the questions. Everyone gets differente numbers in their equations we don't all have the same questions! Would you mind just going over it again

OpenStudy (ttp):

x-3y+4z=-9 2x+y+3z=6 x-4y+3z=-9 solve by elimination method x-3y+4z=-9 multiplie by -2 and will get -2x +6y -8z=18 2x +y +3z = 6 ------------------add term by term 0 7y -5z = 24 2x+y+3z=6 x -4y+3z=-9 multiplie by -2 and will get -2x+8y-6z=18 2x +y +3z=6 -2x+8y-6z=18 --------------- add term by term 0 9y -3z =24 than you have got first 7y-5z=24 multiplie first equation by 3 to eliminate adding the term of z 9y-3z=24 multiplie second equation by -5 21y-15z= 72 -45y+15z=-120 ----------------- adding will get -24y 0 = -48 -24y=-48 y= -48/(-24) = 2 y=2 7y-5z=24 7*2-5z=24 14-5z=24 -5z=10 z= 10/(-5) = -2 z = -2 y = 2 2x+y+3z=6 2x +2 +3(-2)=6 2x +2 -6 =6 2x +4=6 2x= 2 x=2/2 x =1 y = 2 z = -2